OpenStudy (anonymous):

Explain how to find the solution to the system below. Then find the solution(s) to the system. Round answers to the nearest hundredth, if necessary. y = x^2 - x - 12 y = x

OpenStudy (anonymous):

put \[x=x^2-x-12\] and solve for \(x\) dear

OpenStudy (anonymous):

start with \[x^2-2x-12=0\]

OpenStudy (anonymous):

then solve using the quadratic formula or completing the square you pick

OpenStudy (anonymous):

can you hlp me... on it

OpenStudy (anonymous):

lets complete the square because \(-2\) is even \[x^2-2x-12=0\\ x^2-2x=12\\ (x-1)^2=12+1=13\\ x-1=\pm\sqrt{13}\\ x=1\pm\sqrt{13}\]

OpenStudy (anonymous):

thanks